OpenStudy (darr):

Which of the following is a solution of x2 + 5x = −2?

OpenStudy (tkhunny):

Have you considered solving it? Factoring? Completing the square? Quadratic formula? Guess and check?

jhonyy9 (jhonyy9):

<<following>> where ?

OpenStudy (otherworldly):

XD

OpenStudy (shadelily):

First, get them all on one side: \[x ^{2} + 5x + 2 = 0\] Next. I suggest using the quadratic formula \[\frac{ \pm b \sqrt{b ^{2} + 4ac} }{ 2a }\] Which basically means you take the previously given values and assign them letters in order, which gives you: a = 1 b = 5 c = 2 And then just sub in the numbers where you see the letters on the formula. You should get two answers and both of them will be correct. Good luck and hope that helps!